OpenStudy (anonymous):

Find all integers k so that the trinomial 2m^2+km+5 can be factored

OpenStudy (anonymous):

Discriminant of quadratic (in terms of m) must be a perfect square\[k^2-40=n^2\]

OpenStudy (anonymous):

\[(k-n)(k+n)=40\]

OpenStudy (anonymous):

7 and 11 seem to work, they are the only 2